# An average Brun–Titchmarsh theorem and shifted primes with a large prime factor

### v2

@@ 6. The final decomposition
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{802}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in L_{803}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\int_{t_{4}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{804}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{805}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\int_{t_{5}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{2}}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6})\in L_{806}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{t_{6}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}^{2}}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
@@
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{902}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in L_{903}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\int_{t_{4}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{904}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{905}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\int_{t_{5}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{2}}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6})\in L_{906}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{t_{6}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}^{2}}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
$\quad\quad\quad\int_{t_{5}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{2}}\int_{t_{6}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{2}}$
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6},t_{7})\in L_{907}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}-t_{7}}{t_{7}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}t_{7}^{2}}dt_{7}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
$\quad\quad\quad\int_{\kappa}^{\min\left(t_{5},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{2}\right)}\int_{\kappa}^{\min\left(t_{6},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{2}\right)}$
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6},t_{7})\in L_{908}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}-t_{7}}{t_{7}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}t_{7}^{2}}dt_{7}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{909}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in L_{910}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\int_{t_{4}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{911}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa^{\prime}}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa^{\prime}}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{912}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa^{\prime}}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa^{\prime}}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\int_{t_{5}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{2}}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6})\in L_{913}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{t_{6}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}^{2}}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa^{\prime}}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa^{\prime}}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
$\quad\quad\quad\int_{t_{5}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{2}}\int_{t_{6}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{2}}$
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6},t_{7})\in L_{914}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}-t_{7}}{t_{7}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}t_{7}^{2}}dt_{7}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa^{\prime}}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa^{\prime}}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
$\quad\quad\quad\int_{\kappa^{\prime}}^{\min\left(t_{5},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{2}\right)}\int_{\kappa^{\prime}}^{\min\left(t_{6},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{2}\right)}$
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6},t_{7})\in L_{915}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}-t_{7}}{t_{7}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}t_{7}^{2}}dt_{7}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
@@
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{X02}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in L_{X03}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\int_{t_{4}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{X04}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{X05}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\int_{t_{5}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{2}}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6})\in L_{X06}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{t_{6}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}^{2}}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
$\quad\quad\quad\int_{t_{5}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{2}}\int_{t_{6}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{2}}$
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6},t_{7})\in L_{X07}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}-t_{7}}{t_{7}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}t_{7}^{2}}dt_{7}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
$\quad\quad\quad\int_{\kappa}^{\min\left(t_{5},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{2}\right)}\int_{\kappa}^{\min\left(t_{6},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{2}\right)}$
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6},t_{7})\in L_{X08}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}-t_{7}}{t_{7}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}t_{7}^{2}}dt_{7}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$

### v3

@@ 6. The final decomposition
$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{802}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{(t_{1},t_{2},t_{3},t_{4})\in L_{803}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{804}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{805}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6})\in L_{806}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{t_{6}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}^{2}}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
@@
$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{902}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{(t_{1},t_{2},t_{3},t_{4})\in L_{903}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{904}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{905}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6})\in L_{906}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{t_{6}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}^{2}}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6},t_{7})\in L_{907}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}-t_{7}}{t_{7}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}t_{7}^{2}}dt_{7}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6},t_{7})\in L_{908}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}-t_{7}}{t_{7}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}t_{7}^{2}}dt_{7}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{909}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{(t_{1},t_{2},t_{3},t_{4})\in L_{910}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{911}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{912}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6})\in L_{913}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{t_{6}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}^{2}}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6},t_{7})\in L_{914}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}-t_{7}}{t_{7}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}t_{7}^{2}}dt_{7}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6},t_{7})\in L_{915}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}-t_{7}}{t_{7}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}t_{7}^{2}}dt_{7}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
@@
$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{X02}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{(t_{1},t_{2},t_{3},t_{4})\in L_{X03}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{X04}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{X05}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6})\in L_{X06}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{t_{6}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}^{2}}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6},t_{7})\in L_{X07}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}-t_{7}}{t_{7}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}t_{7}^{2}}dt_{7}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6},t_{7})\in L_{X08}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}-t_{7}}{t_{7}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}t_{7}^{2}}dt_{7}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$